# The Feasibility of a Spacecraft Flyby with the Third Interstellar Object 3I/ATLAS from Earth or Mars

### v1

@@ 1 Introduction
Preliminary observations of 3I/ATLAS revealed a weakly active nucleus (D. Jewitt & J. Luu, 2025; M. R. Alarcon et al., 2025; C. O. Chandler et al., 2025) with a reddened reflectance spectrum and an absolute magnitude of $H_{V}\sim 12$ (D. Z. Seligman et al., 2025; B. T. Bolin et al., 2025; C. Opitom et al., 2025) although the nucleus size is uncertain (A. Loeb, 2025). Its visible and near-infrared reflectance is red and featureless, matching the spectral slopes of D-type asteroid and active Centaurus (M. Belyakov et al., 2025; T. Kareta et al., 2025; R. de la Fuente Marcos et al., 2025). 3I/ATLAS’s high excess velocity and Galactic velocity radiant imply an old kinematic age of 3–11 Gyr (A. G. Taylor & D. Z. Seligman, 2025), consistent with an origin from the thick disk of the Milky Way (M. J. Hopkins et al., 2025; S. Kakharov & A. Loeb, 2024).
The first two interstellar objects exhibited remarkably different characteristics. 1I/‘Oumuamua displayed no cometary activity (K. J. Meech et al., 2017; Q.-Z. Ye et al., 2017; D. Jewitt et al., 2017; D. E. Trilling et al., 2018) but its trajectory exhibited a significant nongravitational acceleration (M. Micheli et al., 2018). In stark contrast, 2I/Borisov displayed clear signatures of active coma (D. Jewitt & J. Luu, 2019; A. Fitzsimmons et al., 2019; Q. Ye et al., 2020; A. J. McKay et al., 2020; P. Guzik et al., 2020; M.-T. Hui et al., 2020; Y. Kim et al., 2020; G. Cremonese et al., 2020; B. Yang et al., 2021). 1I/‘Oumuamua and 2I/Borisov had lower excess velocity than 3I/ATLAS, corresponding to younger kinematic ages albeit with large uncertainty (E. Mamajek, 2017; E. Gaidos et al., 2017; F. Feng & H. R. A. Jones, 2018; F. Almeida-Fernandes & H. J. Rocha-Pinto, 2018; T. Hallatt & P. Wiegert, 2020; C.-H. Hsieh et al., 2021).
The discovery of these interstellar objects has ignited significant interest in sending dedicated spacecraft to measure their properties in situ (e.g., D. Seligman & G. Laughlin 2018; K. Moore et al. 2021). Space-based measurements of interstellar object physical and chemical attributes — such as isotopic signatures and volatile compositions — would profoundly enhance our understanding of extrasolar systems at both broad and narrow scales. Compositional indicators such as the H2O content — when combined with the kinematics and galactic history of a given object — can provide insight into planet and star formation rates throughout the galaxy (C. Lintott et al., 2022; M. J. Hopkins et al., 2023, 2025). On the other hand, the carbon-to-oxygen ratio of an interstellar object approximately traces its formation location in a protoplanetary disk (D. Z. Seligman et al., 2022b). Since analogous objects may be responsible for enriching short-period exoplanet atmospheres (D. Z. Seligman et al., 2022a), such data could provide insight into the delivery of volatiles and organics to exoplanets.
@@
The discovery of 1I/‘Oumuamua and 2I/Borisov rapidly spurred development on an interstellar object mission concept. Initial studies (e.g., D. Seligman & G. Laughlin 2018) emphasized the urgency of preparing interception missions with minimal latency but determined that it was feasible to detect and reach future interstellar objects with conventional chemical propulsion. Further studies such as Project Lyra examined the feasibility of direct missions to 1I/‘Oumuamua using near-term technologies, including solar Oberth maneuvers and high-energy launch architectures (A. M. Hein et al., 2019; A. Hibberd et al., 2021; A. Hibberd, 2023). These analyses demonstrated that with optimized trajectories and sufficient launch energies, flyby or rendezvous missions could be viable even years after 1I/‘Oumuamua’s perihelion. Alternative mission design proposals include solar sail architectures capable of prolonged loitering and rapid-response interception of future interstellar objects (A. M. Hein et al., 2019; D. Garber et al., 2022; D. Miller et al., 2022), or a distributed statite cluster for high-$\Delta V$ response scenarios (D. J. Hoover et al., 2022). Evaluations of optical and autonomous navigation performance during high-velocity flybys emphasize the importance of on-board image processing and neural network guidance to maintain targeting accuracy at encounter speeds exceeding 60 km s-1 (D. Mages et al., 2022; B. P. S. Donitz et al., 2023). A complementary analysis focused on the tradeoffs between target detectability, spacecraft storage strategies, and imaging resolution advocated for rendezvous missions to 1I/‘Oumuamua-sized targets (A. Siraj et al., 2023). The “Bridge” New Frontiers-class flyby mission concept would intercept an interstellar object with a dedicated science payload including a guided impactor, ultraviolet/visible/infrared spectrometers, and a high-resolution camera (K. Moore et al., 2021). Additional studies explored broader science cases and mission architectures for interstellar object investigation, including rapid-response trajectories and programmatic feasibility (B. P. Donitz et al., 2021), the astrobiological relevance of interstellar objects as potential carriers of organic material and the propulsion technologies that could enable their exploration (M. Lingam et al., 2023), and the practical constraints and feasibility of near-term rendezvous missions to interstellar objects using advanced but achievable propulsion architectures (D. Landau et al., 2023).
@@
Here, we investigate possible missions that could reasonably intercept 3I/ATLAS in the months following its July 2025 discovery. We specifically explore utilizing existing missions near Earth or Mars and consider favorable minimum distances to 3I/ATLAS. This paper is structured as follows — in Section 2, we detail an algorithm designed to determine direct transfer trajectories and launch dates. In Section 3, we calculate minimum-energy mission trajectories to 3I/ATLAS from Earth achievable in the months following its discovery. In Section 4, we identify additional key launch dates and trajectories for a flyby mission to 3I/ATLAS from Mars. Finally, we conclude in Section 5 with a summary of our results and a final discussion of our suggestions for immediate implementation of an intercept mission with 3I/ATLAS.
@@ 2 Identification of Optimal Trajectories
In this section, we describe the algorithm that we implemented to find minimum-energy trajectories to 3I/ATLAS. This algorithm is based on a methodology to calculate the minimum-energy trajectory between two elliptical orbits (H. Leeghim, 2013), which was later expanded to include hyperbolic trajectories in Section 4 of D. Seligman & G. Laughlin (2018).
The algorithm first solves Kepler’s equations as a function of time given a departure date $\tau_{0}$ that is sampled over a 400-day interval. For the remainder of this paper, a subscript of 0 denotes departure and I/T denote interceptor/target. The flight time is a free parameter in the optimization problem and is denoted by $\Delta\tau$. We obtain the target’s heliocentric position from JPL Horizons, and then simplify the boundary-value problem by enforcing only the interceptor’s universal variable residual. For each candidate departure and flight duration, $\Delta\tau$, we extract the interceptor’s heliocentric state ($\vec{r_{0,I}}$, $\vec{v_{0,I}}$) and compute the target’s position $\vec{r_{T}}$ at arrival. For every pair of both (i) departure date and (ii) arrival date, we solve the minimum-energy single-impulse trajectory by applying the two-body Lambert problem using the universal variable method. With this method, the function $\eta$, which indicates how well the solution fits the Kepler’s equations of motion, may be written as:
$\begin{split}\eta\big{(}\chi,\tau,\vec{r}_{0},\vec{v}_{0}\big{)}=\lVert\vec{r}_{0}\rVert\chi-\sqrt{GM_{\odot}}\Delta\tau\\ +\frac{\vec{r}_{0}\cdot\vec{v}_{0}}{\sqrt{GM_{\odot}}}\,\chi^{2}\,C\bigg{(}\frac{\chi^{2}}{a}\bigg{)}\\ +\left(1-\frac{\lVert\vec{r}_{0}\rVert}{a}\right)\chi^{3}S\bigg{(}\frac{\chi^{2}}{a}\bigg{)}\,.\end{split}$ (1)
In Equation 1, the vectors $\vec{r}_{0}$ and $\vec{v}_{0}$ are the initial position and velocity vectors of the object. The universal variable $\chi$ is defined to be $\chi\equiv\sqrt{a}E$ for elliptical orbits and $\chi\equiv\sqrt{-a}F$ for hyperbolic orbits with semimajor axis $a$ and eccentric/hyperbolic anomaly $E$/ $F$. In Equation 1, $G$ is the gravitational constant, $M_{\odot}$ is the mass of the Sun, and $S(x)$ and $C(x)$ are Stumpff functions, defined as:
@@
For the application of the universal variable formalism to Kepler’s equations, Equations 2 and 3 are typically defined as functions of the variable $x=\chi^{2}/a$.
We calculate the numerical values of the universal variable $\chi$ by finding the roots of the transcendental equation defined by $\eta=0$. Following this step, we use the Lagrange propagators $f(\chi)$ and $g(\chi)$ (and their time derivatives) to advance the interceptor’s state to time $\tau_{0}+\Delta\tau$, since
$\vec{r}(\chi)=f(\chi)\,\vec{r}_{0}+g(\chi)\vec{v}_{0}\,,$ (4)
and
$\vec{v}(\chi)=\dot{f}(\chi)\,\vec{r}_{0}+\dot{g}(\chi)\vec{v}_{0}\,.$ (5)
We refer the reader to the corresponding Equations 25-28 in D. Seligman & G. Laughlin (2018) for definitions of the Lagrange propagators.
Since the arrival position $\vec{r_{T}}$ is taken directly from Horizons, we reformulate the boundary-value problem in terms of the variable $z=\chi^{2}/a$. With this simplification, there is no need to solve the target’s universal variable equation as described in D. Seligman & G. Laughlin (2018).
For a given interceptor state $\vec{r}_{0,I}$, $\vec{v}_{0,I}$, and $\Delta\tau$, the universal variable time-of-flight residual is given by:
$\begin{split}\eta_{I}\big{(}\chi_{I},\tau_{0}+\Delta\tau,\vec{r}_{0,I},\vec{v}_{0}+\vec{v}_{0,I}\big{)}={\lVert\vec{r}_{0,I}\rVert}\chi_{I}-\sqrt{GM_{\odot}}\Delta\tau\\ +\frac{{\vec{r}_{0,I}}\cdot(\vec{v}_{0}+\vec{v}_{0,I})}{\sqrt{GM_{\odot}}}\,\chi_{I}^{2}\,C\bigg{(}\frac{\chi_{I}^{2}}{a}\bigg{)}\\ +\left(1-\frac{\lVert\vec{r}_{0,I}\rVert}{a}\right)\chi_{I}^{3}S\bigg{(}\frac{\chi_{I}^{2}}{a}\bigg{)}\,.\end{split}$ (6)
Rather than propagating both orbits, we require that the interceptor reaches the known target position $\vec{r_{T}}$, (given by Horizons) after a flight time of $\Delta\tau$. To this end, we compute the angle between the target and interceptor, $\Delta\theta$, given by:
We define a geometric coefficient:
$\begin{split}A&=\sin\big{(}\Delta\theta\big{)}\,\sqrt{\frac{\lVert\vec{r}_{T}\rVert\,\lVert\vec{r}_{I}\rVert}{1-\cos\big{(}\Delta\theta\big{)}}}\,,\end{split}$ (8)
which rescales the chord connecting the departure and arrival position vectors so that the universal variable time-of-flight equation properly reflects the transfer geometry. We then introduce the intermediate function:
@@
$\begin{split}F(z)&=\Bigl{(}\frac{y(z)}{C(z)}\Bigr{)}^{3/2}\!S(z)+A\,\sqrt{y(z)}-\sqrt{{GM_{\odot}}}\,\Delta\tau\,,\end{split}$ (10)
Root-finding on Equation 10 enforces both geometric and temporal constrains simultaneously. For a root $z$, the Lambert-Lagrange coefficients are:
@@
By substituting these values into Equations (4, 5), the required arrival position and velocity vectors can be written as:
@@
    Compute the Lagrange coefficient using Equation 11 and propagate the interceptor state with Equations 12 and 13.
@@
It is important to note that these $\Delta V$ values only encapsulate the impulsive burn required to leave a circular low-Earth orbit on an escape trajectory. In other words, they do not include the velocity required to leave the gravitational potential of the departure planet.
@@ 3 Optimal Trajectories from Earth
To determine when a direct launch from Earth could be viable, we use the Earth’s position as the initial launch date for a spacecraft. We then calculate the minimum $\Delta V$ required to reach 3I/ATLAS for a range of possible initial launch dates. The minimum $\Delta V$ required as a function of launch date for a direct flight is shown in Figure 1. The color of each point indicates the flight time for each of the corresponding trajectories.
The minimum $\Delta V$ solution recovered after the discovery of the object technically occurs on 2025 July 1. In other words, the optimal intercept mission would have required immediate launch of a spacecraft contemporaneous with discovery. This optimal trajectory would have required $\Delta V\sim 24.0$ km s-1, with a flight time of 137 days. The details of this trajectory are reported in Table 1. An orbit schematic of this hypothetical minimum-energy flyby mission is shown in Figure 2.
Launch Date Launch Location $V_{x}$ $V_{y}$ $V_{z}$ $\Delta V$ Flight time Arrival Date
[km s-1] [km s-1] [km s-1] [km s-1] [days]
2025 January 10 Earth -5.353 0.592 4.371 6.935 248 2025 September 15
2025 July 1 Earth -8.196 22.536 -0.996 24.001 137 2025 November 15
2025 December 15 Earth 8.424 70.532 -4.082 71.151 198 2026 July 1
2025 March 6 Mars -0.483 -0.082 1.958 2.019 212 2025 October 4
2025 July 1 Mars -1.641 -1.033 2.959 3.538 94 2025 October 3
2025 August 10 Mars -3.291 -1.994 4.8354 6.179 54 2025 October 3
2025 November 10 Mars 14.410 72.571 3.959 74.093 233 2026 July 1
@@
A direct-transfer trajectory from Earth would have been more energetically feasible if 3I/ATLAS had been discovered earlier. Earth departures between January and June would have required $\Delta V\sim 10$ km s-1, with a flight time of approximately 150–200 days. A launch on 2025 January 10 would require only a minimum $\Delta V$ of 6.94 km s-1 with a flight time of 248 days, which may have been within the reach of modern spacecraft. It is worth noting that in a recent and complimentary analyses, A. Hibberd et al. (2025) identified an optimal trajectory from Earth to 3I/ATLAS. They calculated a minimum energy trajectory launched in 2024 July 19 for a rendezvous 2025 October 18. However, that trajectory assumed a continuous acceleration as opposed to a single impulsive burn.
@@
Figure 4: Similar to Figure 2, except for a minimum $\Delta$V mission to 3I/ATLAS post-discovery, sent on 2025 July 1 from Mars. This hypothetical trajectory requires a $\Delta V\sim 4$ km s-1 and a time of flight of 94 days.
@@ 4 Optimal Trajectories from Mars
Mars has been a high-priority planetary science target for multiple space agencies, including NASA and the ESA, for several decades. A typical mission starts with the “prime mission”, which is designed to achieve the originally-funded objectives. Once a prime mission is completed, if the spacecraft can continue operating for a longer period of time, extend missions can be approved with additional science goals. As a spacecraft approaches the end of its operational life (e.g., running out of fuel) an end-of-mission maneuver is designed to safely dispose of the spacecraft. This end-of-mission maneuver can also be used to achieve additional objectives beyond the spacecraft’s initial design capabilities. For example, the Cassini spacecraft was disposed of by sending the spacecraft into Saturn’s atmosphere, taking in situ measurements that could not be accomplished during the active mission and which resulted in major scientific achievements (S. G. Edgington & L. J. Spilker, 2016). Additionally, the Rosetta mission ended its life by descending to the surface of its target comet 67P/Churyumov-Gerasimenko and making observations directly at the surface (A. Accomazzo et al., 2017; M. A. Barucci & M. Fulchignoni, 2017).
@@
In Figure 3, we show the minimum $\Delta V$ required for such a mission. The minimum $\Delta V$ of 2.02 km s-1 occurs for a launch on 2025 March 6 (four months prior to discovery), with flight time of 212 days. The first post-discovery opportunity also occurs contemporaneously with discovery and would have required a $\Delta V$ of 3.54 km s-1 on 2025 July 1. A mission with this trajectory would arrive on 2025 October 3, three months after discovery, with its 94 days in transit. This optimal mission would fly by 3I/ATLAS during its closest approach to Mars. In Figure 4, we show the orbit of this minimum-energy post-discovery mission. For detailed information regarding these trajectories, see Table 1. After discovery, the required $\Delta V$ for flyby climbs gradually, yet remains at $\Delta V\leq 10$ km s -1 through 2025 August 30.
The sharp rise and gap in $\Delta V$ around October 2025 coincides with the closest approach between Mars and 3I/ATLAS, which occurs on 2025 October 3. The optimum trajectories sent prior to the closest approach have flight times such that the flyby occurs close in time to 2025 October 3. In Figure 5, we show the trajectory for a launch on 2025 August 10, 41 days after discovery, which demonstrates one of possible Mars departure scenarios. On the other hand, even the optimal trajectories after the close approach date would require significantly more $\Delta V$ to perform a flyby.
These results suggest that an intercept or flyby from Mars would require less $\Delta V$ than an Earth-based mission. A hypothetical Mars-based mission to 3I/ATLAS would have $\Delta V$ requirements within the performance envelope of modern spacecraft propulsion systems. Therefore, such a Mars-based flyby mission could be feasible.

### v2

@@ 1 Introduction
[…]
@@
The discovery of 1I/‘Oumuamua and 2I/Borisov rapidly spurred development on an interstellar object mission concept. Initial studies (e.g., D. Seligman & G. Laughlin 2018) emphasized the urgency of preparing interception missions with minimal latency but determined that it was feasible to detect and reach future interstellar objects with conventional chemical propulsion. Further studies such as Project Lyra examined the feasibility of direct missions to 1I/‘Oumuamua using near-term technologies, including solar Oberth maneuvers and high-energy launch architectures (A. M. Hein et al., 2019; A. Hibberd et al., 2021; A. Hibberd, 2023). These analyses demonstrated that with optimized trajectories and sufficient launch energies, flyby or rendezvous missions could be viable even years after 1I/‘Oumuamua’s perihelion. Alternative mission design proposals include solar sail architectures capable of prolonged loitering and rapid-response interception of future interstellar objects (A. M. Hein et al., 2019; D. Garber et al., 2022; D. Miller et al., 2022), or a distributed statite cluster for high-$\Delta V$ response scenarios (D. J. Hoover et al., 2022). Evaluations of optical and autonomous navigation performance during high-velocity flybys emphasize the importance of on-board image processing and neural network guidance to maintain targeting accuracy at encounter speeds exceeding 60 km s-1 (D. Mages et al., 2022; B. P. S. Donitz et al., 2023). A complementary analysis focused on the tradeoffs between target detectability, spacecraft storage strategies, and imaging resolution advocated for rendezvous missions to 1I/‘Oumuamua-sized targets (A. Siraj et al., 2023). S. A. Stern et al. (2024) also investigated the feasibility of an intercept mission to an interstellar object, “Interstellar Object Explorer.” The “Bridge” New Frontiers-class flyby mission concept would intercept an interstellar object with a dedicated science payload including a guided impactor, ultraviolet/visible/infrared spectrometers, and a high-resolution camera (K. Moore et al., 2021). Additional studies explored broader science cases and mission architectures for interstellar object investigation, including rapid-response trajectories and programmatic feasibility (B. P. Donitz et al., 2021), the astrobiological relevance of interstellar objects as potential carriers of organic material and the propulsion technologies that could enable their exploration (M. Lingam et al., 2023), and the practical constraints and feasibility of near-term rendezvous missions to interstellar objects using advanced but achievable propulsion architectures (D. Landau et al., 2023).
@@
Here, we investigate possible missions that could reasonably intercept 3I/ATLAS in the months following its July 2025 discovery. We specifically explore utilizing existing missions near Earth or Mars and consider favorable minimum distances to 3I/ATLAS. This paper is structured as follows — in Section 2, we detail an algorithm designed to determine direct transfer trajectories and launch dates. In Section 3, we calculate minimum-energy mission trajectories to 3I/ATLAS from Earth achievable in the months following its discovery. Here, we defined minimum-energy as the trajectory requiring the lowest single‐impulse $\Delta V$. In Section 4, we identify additional key launch dates and trajectories for a flyby mission to 3I/ATLAS from Mars. Finally, we conclude in Section 5 with a summary of our results and a final discussion of our suggestions for immediate implementation of an intercept mission with 3I/ATLAS.
@@ 2 Identification of Optimal Trajectories
We compute minimum-energy transfers using the universal-variable formulation of Lambert’s problem A. A. Sukhanov (1989); R. R. Bate et al. (1971); D. A. Vallado (2001), applied to Earth and a hyperbolic target by following the elliptical-orbit formulation of H. Leeghim (2013) and the hyperbolic-target extension in Section 4 of D. Seligman & G. Laughlin (2018).
We sample departure epochs $\tau_{0}$ over a 400-day window and flight times $\Delta\tau$ across a mission-relevant range. Under a patched-conic approximation, the departure planet’s heliocentric state, we take the heliocentric state $\big(\vec{r}_{0,I},\vec{v}_{h}\big)$ at $\tau_{0}$ and the target position $\vec{r}_{T}$ at $\tau_{0}+\Delta\tau$ are taken from JPL Horizons. Each grid point $(\tau_{0},\Delta\tau)$ is solved as a single universal-variable Lambert problem to obtain the required departure velocity $\vec{v}_{0,I}$ and its associated direct-flight cost $\Delta V$.
We work with the standard universal-variable parameter $z\equiv\frac{\chi^{2}}{a}$ with semimajor axis $a$ and universal variable $\chi$. The transfer geometry is set by the angle between the interceptor position $\vec{r}_{I}\equiv\vec{r}_{0,I}$ and the known target position $\vec{r}_{T}$. We compute the angle between the target and interceptor, $\Delta\theta$, given by:
We define a geometric coefficient:
[…]
which rescales the chord connecting the departure and arrival position vectors so that the universal variable time-of-flight equation properly reflects the transfer geometry.
We then introduce the intermediate function:
Here, $S(z)$ and $C(z)$ denote the Stumpff functions in their conventional definitions (maybe cite).
@@
[…]
Root-finding on Equation 6 enforces both geometric and temporal constrains simultaneously. For a root $z$, the Lambert-Lagrange coefficients are
@@
The required arrival position and velocity vectors can be written as:
@@
    Compute the Lagrange coefficient using Equation 7 and propagate the interceptor state with Equations 8 and 9.
@@
It is important to note that these $\Delta V$ values only encapsulate the hyperbolic-excess component of the impulsive burn required to leave a circular low-Earth orbit on an escape trajectory. In other words, they do not include the velocity needed to climb out of the departure planet’s gravitational potential nor any downstream deep space maneuvers.
@@
The universal variable Lambert solver described above is a well-established solution to a two‐body problem (D. Seligman & G. Laughlin, 2018; H. D. Curtis, 2020). Our novel contribution is not in the solver itself but in its systematic application to the evolving geometry of 3I/ATLAS and subsequent mapping of the resulting minimum $\Delta V$ to the Earth- and Mars-based launchers. This enables the identification of mission-ready intercept windows under realistic departure and arrival constraints.
@@
Figure 2: Mission Design Contour displaying required $\Delta V$ (0–100 km $s^{-1}$) across Earth departure and 3I/ATLAS arrival dates. The Green circle mark indicates the minimum $\Delta V$ on discovery date, 2025 July 1.
@@ 3 Optimal Trajectories from Earth
Launch Date Launch $V_{x}$ $V_{y}$ $V_{z}$ $\Delta V$ Flight time Arrival Date Flyby Speed Phase
Location [km s-1] [km s-1] [km s-1] [km s-1] [days] [km s-1] Angle []
2025 January 10 Earth -5.353 0.592 4.371 6.935 248 2025 September 15 79.96 49.0
2025 July 1 Earth -8.196 22.536 -0.996 24.001 137 2025 November 15 79.73 95.4
2025 December 15 Earth 8.424 70.532 -4.082 71.151 198 2026 July 1 15.84 62.6
2025 March 6 Mars -0.483 -0.082 1.958 2.019 212 2025 October 4 86.19 66.8
2025 July 1 Mars -1.641 -1.033 2.959 3.538 94 2025 October 3 86.43 65.2
2025 August 10 Mars -3.291 -1.994 4.8354 6.179 54 2025 October 3 86.32 64.3
2025 November 10 Mars 14.410 72.571 3.959 74.093 233 2026 July 1 18.88 46.9
@@
To determine when a direct launch from Earth could be viable, we use the Earth’s position as the initial launch date for a spacecraft. We then calculate the minimum $\Delta V$ required to reach 3I/ATLAS for a range of possible initial launch dates. The minimum $\Delta V$ required as a function of launch date for a direct flight is shown in Figure 1. The color of each point indicates the flight time for each of the corresponding trajectories. Figure 2 shows the required $\Delta V$ across Earth departure and 3I/ATLAS arrival date.
The minimum $\Delta V$ solution recovered after the discovery of the object technically occurs on 2025 July 1. In other words, the would have required immediate launch of a spacecraft contemporaneous with discovery. This trajectory would have required $\Delta V\sim 24.0$ km s-1 , with a flight time of 137 days. The details of this trajectory are reported in Table 1. An orbit schematic of this hypothetical minimum-energy flyby mission is shown in Figure 3.
@@
A direct-transfer trajectory from Earth would have been more energetically feasible if 3I/ATLAS had been discovered earlier. Earth departures between January and June would have required $\Delta V\sim 10$ km s-1, with a flight time of approximately 150–200 days. A launch on 2025 January 10 would require only a minimum $\Delta V$ of 6.94 km s-1 with a flight time of 248 days, which may have been within the reach of modern spacecraft. J. P. Sanchez & C. Snodgrass (2025) demonstrated that a complementary deep‐space maneuver (DSM) study for the European Space Agency Comet Interceptor mission (CI) could intercept 3I/ATLAS on 2025 November 7 after an $\sim 825$ days of flight time. This would require departing the Sun-Earth L2 halo orbit on 2023 August 2 with a hyperbolic departure velocity of 2.8 km s-1 and performing a 685 m s-1 midcourse burn. It is worth noting that in a recent and complimentary analyses, A. Hibberd et al. (2025) identified an optimal trajectory from Earth to 3I/ATLAS. They calculated a minimum energy trajectory launched in 2024 July 19 for a rendezvous 2025 October 18. That trajectory assumes a single impulsive $\Delta V$ burn
@@
Figure 5: Similar to Figure 2, but from Mars in this case.
[…]
@@ 4 Optimal Trajectories from Mars
Mars has been a high-priority planetary science target for multiple space agencies, including NASA and the ESA, for several decades. A typical mission starts with the “prime mission”, which is designed to achieve the originally-funded objectives. Once a prime mission is completed, if the spacecraft can continue operating for a longer period of time, extended missions can be approved with additional science goals. As a spacecraft approaches the end of its operational life (e.g., running out of fuel) an end-of-mission maneuver is designed to safely dispose of the spacecraft. This end-of-mission maneuver can also be used to achieve additional objectives beyond the spacecraft’s initial design capabilities. For example, the Cassini spacecraft was disposed of by sending the spacecraft into Saturn’s atmosphere, taking in situ measurements that could not be accomplished during the active mission and which resulted in major scientific achievements (S. G. Edgington & L. J. Spilker, 2016). Additionally, the Rosetta mission ended its life by descending to the surface of its target comet 67P/Churyumov-Gerasimenko and making observations directly at the surface (A. Accomazzo et al., 2017; M. A. Barucci & M. Fulchignoni, 2017).
@@
In Figure 4, we show the minimum $\Delta V$ required for such a mission. The minimum $\Delta V$ of 2.02 km s-1 occurs for a launch on 2025 March 6 (four months prior to discovery), with flight time of 212 days. The first post-discovery opportunity also occurs contemporaneously with discovery and would have required a $\Delta V$ of 3.54 km s-1 on 2025 July 1. A mission with this trajectory would arrive on 2025 October 3, three months after discovery, with its 94 days in transit. This optimal mission would fly by 3I/ATLAS during its closest approach to Mars. In Figure 5, we show the required $\Delta V$ as a function of the Mars departure date and 3I/ATLAS arrival date. In Figure 6, we show the orbit of this minimum-energy post-discovery mission. For detailed information regarding these trajectories, see Table 1. After discovery, the required $\Delta V$ for flyby climbs gradually, yet remains at $\Delta V\leq 10$ km s -1 through 2025 August 30.
The sharp rise and gap in $\Delta V$ around October 2025 coincides with the closest approach between Mars and 3I/ATLAS, which occurs on 2025 October 3.
The trajectories sent prior to the closest approach have flight times such that the flyby occurs close in time to 2025 October 3. In Figure 7, we show the trajectory for a launch on 2025 August 10, 41 days after discovery, which demonstrates one of possible Mars departure scenarios. On the other hand, even the trajectories after the close approach date would require significantly more $\Delta V$ to perform a flyby.
These results suggest that an intercept or flyby from Mars would require less $\Delta V$ than an Earth-based mission. This advantage is due primarily to orbital phasing. At the perihelion, Mars is closely aligned with 3I/ATLAS orbital longitude, while Earth is nearly $180^{\circ}$ out of phase, making an Earth‐based intercept far more $\Delta V$ intensive. A hypothetical Mars-based mission to 3I/ATLAS would have $\Delta V$ requirements within the performance envelope of modern spacecraft propulsion systems. Therefore, such a Mars-based flyby mission could be feasible.